A 6 m stick of mass 60 kg is rested upon a wall. At what distance should the bottom of the stick be so that it tends to slip? Coeff of friction between stick and floor is 0.3

Grade:11

1 Answers

Vikas TU
7 years ago
Let at x distance fro the bottom it would teds to slip.
Hence the torque to its cetre of mass = 0.
Therfore,
(60x/l)g x (x/2) = 60(l-x)^2/2l
on solving we get,
x = 6/(1+ root(g)) = > 1.452 meter approx.